# Deep Learning Reduced Order Modelling on Parametric and Data driven domains

### v1

@@ 3 Our method
: Neural networks $\Psi_{D}^{S},\Psi_{E}^{D}$, and $\Phi_{S}$, values of parameters and/or domains for which we want to evaluate solution $(\bm{\lambda}_{i},\bm{c}_{i}),i=1,\ldots M$, mean and standard deviation of training parameters $\bm{f}_{i}$ $\mu_{p},\sigma_{p}$, mean and standard deviation of solutions $\mu_{s},\sigma_{s}$
@@
Concatenate $f_{i}=(\bm{\lambda}_{i},\bm{d}_{i}),i=1,\ldots M$
@@
Return solution $\sigma_{s}\bm{\hat{u}}+\mu_{s}$
Experiments and Results
Through this section, we will use relative error to evaluate our approach. The relative error $\epsilon$ for sample $(\bm{\lambda},\bm{u}_{h},\bm{d}_{h})$ is:
$\epsilon(\bm{u}_{h},\hat{\bm{u}_{h}})=\frac{\sqrt{\sum_{i,j}(u^{i,j}_{h}d^{i,j}_{h}-\hat{u}^{i,j}_{h}d^{i,j}_{h})^{2}}}{\sqrt{\sum_{i,j}(u^{i,j}_{h})^{2}d^{i,j}_{h}}},$
where $\bm{\lambda}$ is the parameter vector, $\bm{u}_{h}$ solution discretized on a mesh with $n\times m$ points, $\bm{\hat{u_{h}}}$ is DL-ROM prediction and $\bm{d}_{h}$ is an approximation of characteristic functions of the domains on that same mesh.
Relative error at i-th row and j-th column is:
$\epsilon_{i,j}(u_{k},\hat{u}_{h})=\frac{u_{h}^{i,j}d_{h}^{i,j}-\hat{u}^{i,j}_{h}d^{i,j}_{h}}{\sqrt{\sum_{k,l}(u^{k,l}_{h})^{2}d^{k,l}_{h}}},$
@@
We use FEM implementation in FreeFEM [43] to obtain solutions and Pytorch [44] for defining and training neural networks. All solutions are projected on a rectangular equidistant mesh, in order to use CNN. In all examples, Leaky ReLU is used as an activation function after convolution. We use batch normalization [45] after each layer, except for the last layers of the encoder and decoder. In all examples, the number of channels per layer in the decoder mirrors the encoder architecture. Upsampling operations are used only in some layers, such that the resolution of outputs in the decoder layers mirrors the resolution of outputs in the encoder layers. For all networks, Adam optimizer [46] with One Cycle Policy scheduling [47] is used that adjusts the learning rate at each training step, with a maximum of 1e-3 for the encoder-decoder network.
All training, test, and validation sets used have $1000$ samples. The batch size for the autoencoder is $50$. Autoencoder network hyperparameters were chosen by trial and error with respect to autoencoder reconstruction, where we searched on latent dimensions $15,20,30$ and the number of channels for each encoder layer (the decoder mirrors the encoder). Tried hyperparameters are shown in Table 2. Autoencoders with more channels per layer had better training and test accuracy, so we picked the best two for further analysis.
Latent dimension $[15,20,30]$
Nbr. of kernels in encoder $[4,4,8,8,16,16,16]$, $[4,8,16,32,32,64,64]$ , $[8,16,16,32,64,64,64]$
Learning rate 1e-3
@@
Table 2: Autoencoder grid search hyperparameters. We trained 9 models with varying channel sizes and latent dimension sizes. The number of channels in decoder layers mirrors the encoder number of channels per layer. We have chosen two autoencoders with the smallest reconstruction errors on the test set, both with $8,16,16,32,64,64,64,64$ channels in encoders and dimensions $20$ and $30$ for further analysis.
We have not encountered any overfitting problems by increasing autoencoder capacity. However, neural network $\Phi_{S}$ easily overfits. Since $\Phi_{S}$ is usually small compared to autoencoder and trains fast, we employed grid search to choose the best hyperparameters for $\Phi_{S}$. Hyperparameters are shown in Table 3. The grid search was done for two ROM representations, obtained from solution autoencoders with $8,16,16,32,64,64,64$ channels per layer and latent dimensions equal to $15$ and $30$, respectively. Domain features are obtained from autoencoders trained on $50000$ domains with $32,32,64,64,64,64,64,64$ channels per layer in each layer and strides $1,1,2,2,2,1,1,1$ per layer respectively. For each model, the One Cycle Policy scheduler with Adam optimizer was used. We have trained two different domain autoencoders, one with latent dimension $15$ and one with latent dimension $30$. We choose the best hyperparameters with respect to the mean-squared error between the prediction of $\Phi_{S}$ and ROM representation.
@@
Table 3: Hyperparameters grid search for $\Phi$ is performed, which results in total $4200$ models.
@@ 3.1 Advection-diffusion on different domains
Advection-diffusion equation on parametric domains:
$\begin{cases}-\Delta u+0.1\mathbf{c}\cdot\nabla u=1&\text{on $\langle 0,1\rangle^{2}$ without elipse}\\ u=1&\text{ on edge of ellipse}\\ u=0&\text{for $y=0$ and $y=1$}\\ \frac{\partial u}{\partial\nu}=\beta&\text{ for $x=0$}\\ \frac{\partial u}{\partial\nu}=-\beta&\text{ for $x=1$}\end{cases}$ (4)
@@
  • center of ellipse $(x_{0},y_{0})\in[0.1,0.9]^{2}$
  • angle of major axis of ellipse and x-axis $\alpha\in[0,\pi]$
  • $\beta\in[1,10]$
  • half-axes for ellipse are then taken randomly from interval depending on the center and angle of the ellipse
In total, there are $7$ parameters.
@@
In this example, we know exact geometric parametrization, so we can obtain DL-ROM performance with calculated and exact parametrization and directly compare those two. For each domain parametrization, we have trained a fully-connected neural network $\Phi_{S}$ with hyperparameters determined by grid search.
@@
Comparison of vanilla DL-ROM and our approach is shown in Table 4. We compare DL-ROM performance when exact ellipse parameters are used with domain parametrization calculated with the autoencoder. In all cases, models with calculated ellipse parameters perform better on the test set than models with exact ellipse parameters. In fact, we report that more than $1300$ models out of $4200$ trained in grid search, performed better on parameters of domain obtained autoencoder than the best model trained on exact domain parameters, for each combination. In this case, it can be seen that vanilla DL-ROM is more sensitive to ROM dimension, than our approach. Results are shown on Figure 6.
@@
test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 exact ellipse parameters only 0.0554552 0.0554552 0.0605261 0.0605261 exact and calculated elipse parameters 0.0454250 0.0425800 0.0458577 0.0438216 calculated elipse parameters 0.0465742 0.0452926 0.0451946 0.0449190
@@
Table 4: Comparison of DL-ROM and DL-DA ROM performance: Difference of performance on relative test errors depending on ROM encoding dimension and domain encoding dimension. We differentiate between cases where geometric parameters are exactly known (1st row), when we add domain-encoded features to exact geometric features (2nd row), and when we use only geometric parameters obtained by autoencoder (3rd row). We can see that using autoencoder domain parametrization is beneficial for model performance. We report on the best model for each ROM encoding dimension and domain encoding dimension. Models trained on calculated domain parameters perform better than models trained using only exact domain parameters.
@@
Figure 6: Top row: Exact solution, left, network prediction for a model trained on exact domain parameters (center), network prediction for a model trained on approximated domain parameters. Bottom row: relative error for a model trained on exact domain parameters (left), relative error for a model trained on approximated domain parameters (middle), relative error for a model trained on both exact and calculated domain parameters (right). Results are shown for ellipse with center $(0.76,0.38)$, half-axes $a=0.14,b=0.11$ and angle $\alpha=1.51$, $\beta=6.92$, $\mu=0.29$. All models were trained on $1000$ samples.
@@ 3.2 Domains with varying number of parameters
We experiment again on the advection-diffusion equation on parametric domains. In this case, domains can depend on a different number of parameters. Such a domain may be difficult to parametrize by classical techniques. We are investigating domains that are squares with $1$ to $4$ circular holes. Equations are:
$\begin{cases}-\Delta u+0.1\mathbf{c}\cdot\nabla u=1&\text{on $\langle 0,1\rangle^{2}$ whitout circles}\\ u=2&\text{ on edge of circles}\\ u=0&\text{ for $y=0$ and $y=1$}\\ \frac{\partial u}{\partial\nu}=\beta&\text{ for $x=0$}\\ \frac{\partial u}{\partial\nu}=-\beta&\text{ for $x=1$}\end{cases}$ (5)
@@
  • centers of holes $(x_{i},y_{i})$, $i=1,\ldots m$.
  • circles radii $r_{i}$, $i=1,\ldots m$.
  • $\beta\in[1,10]$
@@
To generate domains, first, we randomly choose the number of circles $m$. Then we randomly choose the centers of circles so the distance between the centers and the square boundary is at least $0.1$. Finally, radii are chosen randomly so that the distance between all circles is at least $0.1$ and $r\in[0.05,0.2]$. This procedure ensures that circles do not overlap and that all circles are of reasonable size. We have chosen the same encoder-decoder architectures as in the previous example and after training solution autoencoders, a grid search was performed to obtain $\Phi_{S}$ hyperparameters. Results are shown in Table 5 and Figure 7. A larger ROM dimension mostly does not affect DL-ROM performance, while a smaller domain parametrization encoding dimension is beneficial in this case.
test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 calculated domain parameters 0.0239904 0.0264217 0.0209590 0.0267886
@@
Table 5: Difference of performance on relative test errors depending on ROM encoding dimension and domain encoding dimension. In this case, domains have varying numbers of parameters. We use only geometric parameters obtained by the autoencoder. We report on the best model for each ROM encoding dimension and domain encoding dimension.
@@
Figure 7: Exact solutions (left), model prediction (middle), relative error (right) for model trained on $1000$ random samples for parameters $\beta=0.977$, $\mu=0.842$, with centers $(0.23,0.73),(0.57,0.22),(0.25,0.37),(0.79,0.79)$ and radii $0.11,0.10,0.14,0.10$ respectively.
Furthermore, we test how sensitive DL-DA ROM DL-ROM is on small perturbations of the circle boundary. The sensitivity of this model is especially important if we train on synthetically generated data and test only on possibly imprecisely measured domains and parameters. Instead of testing the trained model on a domain with circular holes, we test the model on predicting the correct solution when the domain is a unit square without ellipses. The idea is that these ellipses model slightly deformed circles. The correct solution is again generated using the finite element method. We tested on solutions with parameters $\beta=1.19617,\theta=1.71657$ with $4$ holes, with centers $(0.66555,0.35233),(0.56892,0.71108),(0.25424,0.21219),(0.22812,0.58991)$ and radii of non-deformed circles $0.11191,0.12318,0.10707,0.10653$ respectively. A comparison between FEM and our DL-ROM is shown in Figure 8. We tested how sensitive our model is to perturbations of the boundary. In these cases, instead of circular holes, we have elliptic holes. The relative error for increasing deformation is given in Table 6.
Ratio between small and large half-axis two half-axes in % $\epsilon$
1 0.0101
0.95 0.0120
0.9 0.0142
0.8 0.0166
Table 6: Relative errors for the same $\beta$, $\theta$ and ellipse centers as in 8 ($\beta=1.19617$, $\theta=1.71657$). Ellipses have centers $(0.66555,0.35233),(0.56892,0.71108),(0.25424,0.21219),(0.22812,0.58991)$, radii of non-deformed circles $0.11191,0.12318,0.10707,0.10653$ respectively and varying rations between small and large half-axis. We can see an increase in relative error with higher deformation of the domain boundary, but the increase is not rapid.

### v2

@@ 3 Our method
: Neural networks $\Psi_{D}^{S},\Psi_{E}^{D}$, and $\Phi_{S}$, values of parameters and/or domains $\bm{d}$ for which we want to evaluate solution $(\bm{\lambda}_{i},\bm{c}_{i}),i=1,\ldots M$, mean and standard deviation of training parameters $\bm{f}_{i}$ $\bm{\mu}_{p},\bm{\sigma}_{p}$, mean and standard deviation of solutions $\mu_{s},\sigma_{s}$
@@
Concatenate $\bm{f}_{i}=(\bm{\lambda}_{i},\bm{d}_{i}),i=1,\ldots M$
@@
Return solution $(\sigma_{s}\bm{\hat{u}}+\mu_{s})\odot d$
4 Experiments and Results
Throughout this section, we will use relative error to evaluate our approach. The relative error $\epsilon$ for a sample $(\bm{\lambda},\bm{u}_{h},\bm{d}_{h})$ is:
$\epsilon(\bm{u}_{h},\hat{\bm{u}_{h}})=\frac{\sqrt{\sum_{i,j}(u^{i,j}_{h}-\hat{u}^{i,j}_{h})^{2}d^{i,j}_{h}}}{\sqrt{\sum_{i,j}(u^{i,j}_{h})^{2}d^{i,j}_{h}}},$
where $\bm{\lambda}$ is the parameter vector, $\bm{u}_{h}$ solution discretized on a mesh with $n\times m$ points, $\bm{\hat{u_{h}}}$ is DL-ROM prediction and $\bm{d}_{h}$ is an approximation of characteristic functions of domains on that same mesh.
For visualization purposes, we also define a pointwise relative error at position $(i,j)$:
$\epsilon_{i,j}(u_{k},\hat{u}_{h})=\frac{(u_{h}^{i,j}-\hat{u}^{i,j}_{h})d^{i,j}_{h}}{\sqrt{\sum_{k,l}(u^{k,l}_{h})^{2}d^{k,l}_{h}}},$
@@ 4 Experiments and Results
We use a FreeFEM++ software [42] to generate finite element solutions and PyTorch [43] to implement and train all neural networks. All solutions are projected onto a uniform rectangular grid to enable the use of CNNs. For the advection–diffusion examples, the elliptical-hole and circular-hole cases are discretized on a $128\times 128$ grid, while the Navier–Stokes case uses a $256\times 512$ grid. Interpolation relative error is calculated as L2 relative error of the data on interpolated mesh projected to original mesh and original data and is approximately $0.06\%$ for every example. In all experiments, we use Leaky ReLU activations in convolutional layers and apply batch normalization [44] after each layer, except for the final layers of the encoder and decoder. The decoder mirrors the encoder in architecture and number of channels. Upsampling layers are used in selected layers to ensure output resolutions match those of the corresponding encoder layers. All networks are trained using Adam optimizer [45] with the One Cycle Policy learning rate scheduler [46], with a maximum learning rate of $1\times 10^{-3}$ for autoencoder networks.
All training, test, and validation sets used have $1000$ samples. The batch size for the autoencoder is $50$. Autoencoder network hyperparameters are chosen by trial and error with respect to autoencoder reconstruction, where we searched on latent dimensions $15$, and $30$ and the number of convolutional channels for each encoder layer (the decoder mirrors the encoder). The hyperparameters tested are shown in Table 2, are identical across all examples. Autoencoders with more channels per layer have better training and test accuracy, so we picked the best two for further analysis.
Latent dimension $15,30$
Nbr. of channels in each
encoder layer
$[4,4,8,8,16,16,16]$,
$[4,8,16,32,32,64,64]$ ,
$[8,16,16,32,64,64,64]$
Strides
$[1,1,2,2,2,1,1]$
Maximum learning rate 1e-3
@@
Table 2: Autoencoder grid search hyperparameters. We trained 6 models with varying channel sizes and latent dimension sizes. The number of channels in decoder layers mirrors the encoder number of channels per layer. We have chosen two autoencoders with the smallest mean relative reconstruction errors on the validation set, both with $8,16,16,32,64,64,64$ channels in encoders and dimensions $15$ and $30$ for further analysis.
Since $\Phi_{S}$ is usually small compared to autoencoder and fast to train, we employ grid search to choose the best hyperparameters for $\Phi_{S}$. Hyperparameters are shown in Table 3. The grid search was done for two ROM representations, obtained from solution autoencoders with $8,16,16,32,64,64,64$ channels per layer and latent dimensions equal to $15$ and $30$, respectively. Domain features are obtained from autoencoders trained on $9000$ domains with $8,16,16,32,64,64,64$ channels per layer in each layer and strides $1,1,2,2,2,1,1$ per layer, respectively. For each model, the One Cycle Policy scheduler with Adam optimizer was used. We trained two different domain autoencoders, one with latent dimension $20$ and one with latent dimension $30$. We choose the best hyperparameters with respect to the mean-squared error between the prediction of $\Phi_{S}$ and ROM representation on validation set.
@@
Table 3: Hyperparameters grid search for $\Phi$ is performed, which results in $4200$ models in total.
@@ 4.1 Advection-diffusion on different domains
In this section, we evaluate the accuracy and generalization capabilities of our proposed DL-DA-ROM method on a parameterized advection-diffusion problem. The goal is to compare the performance of models using exact domain parameters (e.g., ellipse descriptors) versus those using domain encodings learned by the autoencoder. This setting allows for a controlled comparison, as both the physical and geometric parameters are known. By analyzing prediction accuracy across different latent dimensions, we assess the quality of the learned domain representations and their ability to replace traditional parameterizations. Precisely, we consider the following advection-diffusion equation on parametric domains:
[…]
@@
  • center of the ellipse $(x_{0},y_{0})\in[0.15,0.85]^{2}$
  • angle between major axis of the ellipse and the x-axis $\alpha\in[0,\pi]$
  • boundary flux $\beta\in[1,10]$
  • half-axes $a,b\in[0.05,0.25]$
In total, there are $7$ parameters: two PDE parameters—transport angle $\phi$ and boundary flux $\beta$—and five geometric parameters defining the ellipse: center coordinates $(x_{0},y_{0})$, half-axes $a$ and $b$, and orientation angle $\alpha$. We assume that the ellipse has to be at least 0.01 units away from each of the boundaries to avoid geometric artifacts.
@@
Since the domain is generated from known ellipse parameters, we can directly compare the performance of our model using either exact geometric parameters or those inferred by the domain autoencoder. For each type of domain parametrization, we train a fully connected neural network $\Phi_{S}$ with hyperparameters selected via grid search.
@@
A comparison of the vanilla DL-ROM and our approach is shown in Table 4. We compare DL-ROM performance when exact ellipse parameters are used with domain parametrization calculated with the autoencoder. Performance of the best models is shown in Figure 6, evaluated on one of the examples from the test set.
@@
test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 exact ellipse parameters only 0.0348 0.0348 0.0383 0.0383 calculated elipse parameters 0.0339 0.0361 0.0310 0.03975
@@
Table 4: Comparison of DL-ROM and DL-DA ROM performance: Difference of the performance on relative test errors depending on the ROM encoding dimension and domain encoding dimension. We differentiate between cases where geometric parameters are exactly known (1st row) and when we use only geometric parameters obtained by autoencoder (2nd row). We can see that using the autoencoder domain parametrization does not degrade model performance. We report on the best model for each ROM encoding dimension and domain encoding dimension. Models trained on calculated domain parameters perform similar as models trained using only exact domain parameters. Even though results are similar, larger fully-connected neural network is needed in case of calculated domain parameters (4 layers with 2048 neurons vs 4 layers with 512 neurons).
@@
Figure 6: Comparison of predictions and relative errors for the DL-ROM with ROM encoding dimension $30$ and domain encoding dimension $20$. From left to right: Exact solution, network prediction for a model trained on exact domain parameters, network prediction for a model trained on approximated domain parameters, relative error for the model trained on exact domain parameters, and relative error for the model trained on approximated domain parameters. Results correspond to the ellipse with center (0.815, 0.456)(0.815,\,0.456), half-axes $a=0.194$, $b=0.110$, transport angle $\phi=2.497$, and boundary flux $\beta=6.884$. All models were trained on $1000$ samples.
In order to compare the exact and autoencoder domain parametrization, we investigate error distributions for models with the best performance according to Table 4. Histograms of relative errors are in Figure 7. The results indicate that domain encoding improves model robustness, as reflected by a reduced presence of outliers.
Figure 7: Histograms of relative errors on test set for model with ROM encoding dimension 30, trained on the exact elipse parameters (blue) and model with ROM encoding dimension 30, trained on 20-dimensional domain encoding (orange). When using exact parametrization, $15$ samples have the relative error larger than $0.1$ and $213$ larger than $0.05$. In case of using calculated domain parametrization, $15$ samples have the relative error larger than $0.1$ and $121$ larger than $0.05$. There are more error outliers when using exact ellipse parametrization.
@@ 4.2 Domains with varying number of parameters
We experiment again with the advection-diffusion equation on parametric domains. In this setting, the domains vary in complexity by depending on a different number of parameters, which makes them challenging to parametrize using classical techniques. Specifically, we consider square domains that contain between one and four circular holes. More precisely, we consider the following boundary value problem:
[…]
@@
  • centers of holes $(x_{i},y_{i})$, $i=1,\ldots m$
  • circles’ radii $r_{i}$, $i=1,\ldots m$
  • boundary flux $\beta\in[1,10]$
In total, we again have two PDE parameters: transport angle $\phi$ and boundary flux $\beta$, but there are between three and twelve geometric parameters, namely, the radius and center coordinates for each circular hole.
@@
To generate domains, first, we randomly choose the number of circles $m$. Then we randomly choose the centers of the circles so that the distance between the centers and the outer boundary is at least $0.1$. Finally, radii are chosen randomly so that the distance between all circles is at least $0.1$ and $r\in[0.05,0.2]$. This procedure ensures that circles do not overlap and that all circles are of reasonable size. We have chosen the same encoder-decoder architectures as in the previous example, and after training solution autoencoders, a grid search was performed to obtain $\Phi_{S}$ hyperparameters. Table 5 shows the average relative error on the test set for different combinations of domain and solution latent dimensions, while Figure 8 shows an example of a solution corresponding to a geometric configuration from the test set.
test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 calculated domain parameters 0.0252 0.0242 0.0244 0.0234
@@
Table 5: Test relative errors obtain usign different ROM and the domain encoding dimensions. In this case, domains have varying numbers of parameters. We use only geometric parameters obtained by the autoencoder. We report on the best model for each ROM encoding dimension and domain encoding dimension. A smaller domain encoding dimension is beneficial in this case.
@@
Figure 8: Exact solutions (left), model prediction (middle), relative error (right) for model trained on $1000$ random samples for parameters $\beta=1.671$, $\phi=0.182$, with centers $(0.322,0.379)$ and $(0.374,0.736)$, and radii $0.130$ and $0.124$, respectively.
Robustness to Domain Deformations. To test the robustness of the model, we evaluated DL-ROMs trained on domains with circular holes using test domains in which the circles were slightly deformed into ellipses. Specifically, each circle was transformed by reducing its minor semiaxis to (1p)(1-p) times the major semiaxis, where $p$ is a deformation factor. To introduce variability, even-numbered ellipses were aligned along the $x$-axis, and odd-numbered ones were aligned along the $y$-axis. This setup simulates realistic domain perturbations, such as geometric noise or measurement uncertainty.
The trained models were tested on these deformed domains, and their predictions were compared with finite element solutions. Table 6 reports the average test errors for increasing levels of deformation. As expected, the error increases as the deformation grows. However, the model does not break down and continues to produce reasonable approximations even when tested on domains with types of geometries not seen during training. This suggests that the method generalizes well to moderate domain perturbations.
Figure 9 shows one such out-of-distribution example. The relative error is highest near the boundaries of the ellipses, where the geometry most significantly deviates from the training distribution. Still, the model captures the global solution behavior accurately. The histogram of test errors in Figure 10 confirms that even under moderate distribution shifts, no significant outliers appear, indicating that the predictions remain stable and that the method is robust.
test error, ROM encoding dimension 15, domain encoding dimension 20 test error, ROM encoding dimension 15, domain encoding dimension 30 test error, ROM encoding dimension 30, domain encoding dimension 20 test error, ROM encoding dimension 30, domain encoding dimension 30 $p=0$ 0.0252 0.0242 0.0244 0.0234 $p=0.05$ 0.0260 0.0252 0.0252 0.0254 $p=0.1$ 0.0280 0.0275 0.0278 0.0294 $p=0.15$ 0.0309 0.0310 0.0319 0.0356
Table 6: Relative test errors for DL-ROMs trained on domains with circular holes ($p=0$), evaluated on domains with increasingly deformed elliptical holes. Here, 1p1-p denotes the ratio between the smaller and larger semi-axis of the ellipse. We report results for the best model for each ROM and domain encoding dimension. Performance degrades gradually with increasing deformation, but remains acceptable, demonstrating robustness to moderate boundary changes.